OpenStudy (freckles):

Solve for a(x+y) given that a is negative, (x+y) is positive, and 2ax^2+4axy+2ay^2-ax-ay=1

OpenStudy (freckles):

@satellite73 you want this one? :p

OpenStudy (anonymous):

no

OpenStudy (freckles):

Aw, it isn't that bad.

OpenStudy (anonymous):

have you noticed how the damned chat boxes sometimes cover up your post button ?

OpenStudy (freckles):

Yes, it is very annoying. I have to refresh whenever I want to post something.

OpenStudy (anonymous):

who should we tell?

OpenStudy (freckles):

I don't know. I gave up with that quest.

ganeshie8 (ganeshie8):

there is a keyboard shortcut for Posting

OpenStudy (anonymous):

really!? all this time and i didn't know it? btw i think this problem might be easy

OpenStudy (freckles):

I was trying to hide the fact it was easy.

ganeshie8 (ganeshie8):

press <Tab> followed by <Enter>

OpenStudy (freckles):

I tried to find a mask for it that would make the problem scary.

OpenStudy (freckles):

or appear scary

OpenStudy (freckles):

i tried the tabbing thing let me try again

OpenStudy (freckles):

omg @ganeshie8 genius

OpenStudy (anonymous):

hi

OpenStudy (freckles):

I tabbed too much

OpenStudy (freckles):

in the past

OpenStudy (freckles):

that was what my problem was

OpenStudy (anonymous):

holy unicorn freckles

OpenStudy (anonymous):

yes, i drank a lot of that diet coca cola product too

OpenStudy (anonymous):

used it to chase my blotter

OpenStudy (freckles):

I don't like aspartame

OpenStudy (freckles):

wait is that how you spell it

OpenStudy (anonymous):

lsd?

OpenStudy (freckles):

ever since they told me they put that stuff in dead people it grossed me out

OpenStudy (anonymous):

eeew

OpenStudy (anonymous):

if you tab and enter all kinds of different stuff happens

OpenStudy (anonymous):

this is great i am going to use it all the time

OpenStudy (freckles):

This problem?

OpenStudy (freckles):

not apartame?

OpenStudy (anonymous):

aspartame i assumed you were making a drug reference when you said you tabbed all the time

OpenStudy (anonymous):

tab enter tab enter tab enter last time i had this eye opener was when someone showed me to to find the latex so i could copy and paste

OpenStudy (freckles):

no i meant when I tried to do the whole tabbing thing gan mentioned i would tab too much at one time and that is why it never worked for me

OpenStudy (freckles):

i assumed i had to hit tab 4 times to get from equation to draw to attach file to post

OpenStudy (anonymous):

ooh i see i get \(-\frac{1}{2}\)?

zepdrix (zepdrix):

The chat box blocking the post button >> This is really only a temporary solution, but it always works. I post a short little comment if I can, then I do any other answering from my profile, questions answered. I don't run into the problem there.

OpenStudy (anonymous):

?

OpenStudy (freckles):

-1 is what i got

OpenStudy (anonymous):

you sure?

OpenStudy (freckles):

I think let me do the problem again

OpenStudy (anonymous):

i like my answer

OpenStudy (anonymous):

tab enter lolololo

OpenStudy (freckles):

\[2a(x^2+2xy+y^2)-a(x+y)-1=0 \\ 2a(x+y)^2-a(x+y)-1=0 \\ a(x+y)=\frac{1 \pm \sqrt{1-4(2)(-1)}}{2(2)}=\frac{ 1 \pm \sqrt{9}}{4}=\frac{1+3}{4} \text{ or } \frac{1-3}{4}\] a(x+y)=-2/4=-1/2

OpenStudy (freckles):

i forgot the 2 on bottom

OpenStudy (freckles):

I can't even make my problems anymore

OpenStudy (freckles):

you win @satellite73

OpenStudy (anonymous):

lol i just hit tab, enter, and reported you!!

OpenStudy (amistre64):

i tend to do tab-space

OpenStudy (amistre64):

but sometimes that takes me to the ask a question box

OpenStudy (freckles):

I want to tab report you let me see if i can figure this out

OpenStudy (freckles):

How many times you press tab? like 1000?

OpenStudy (anonymous):

no once! this report box came up i duly filled it in, claiming that you were an evil twin, and sent it

OpenStudy (anonymous):

tab space works too

OpenStudy (freckles):

"i tabbed, entered and got this!~"

OpenStudy (anonymous):

oh wow

OpenStudy (anonymous):

every damn thing works, who knew

OpenStudy (freckles):

I wanted to be an evil twin.

OpenStudy (anonymous):

keep tabbing, you will be one

OpenStudy (anonymous):

tab space simultaneous take you to feedback, whatever that is

OpenStudy (anonymous):

i see it worked !

myininaya (myininaya):

@ganeshie8 you shouldn't have game @satellite73 a magic wand. No telling what he will do with it.

myininaya (myininaya):

gave*

OpenStudy (amistre64):

tabbing in the chat box just sends you to the chat kernel lol

OpenStudy (anonymous):

i got a twitter feed!!

OpenStudy (anonymous):

tab entered with no text

OpenStudy (anonymous):

now i get nothing think i broke it

myininaya (myininaya):

Maybe you tabbed into a virus

OpenStudy (anonymous):

nah i am running linux, not viruses

OpenStudy (amistre64):

the ebony virus maybe .... the ivory virus is just to ticklish

OpenStudy (anonymous):

@amistre64 no ebola jokes

OpenStudy (anonymous):

ooh you beat me to it, i knew it was going to be an ebola joke

OpenStudy (amistre64):

::murmur::

OpenStudy (anonymous):

ivory virus too ticklish? ho ho ho i hear spike jones is making a comeback

OpenStudy (amistre64):

nah, dead is dead ... theres no hope for spike

myininaya (myininaya):

Oh no. I got a warning.

OpenStudy (amistre64):

i got a notif that myin got a warning ...

OpenStudy (anonymous):

https://www.youtube.com/watch?v=BavRrRNvz8g

OpenStudy (anonymous):

a notif motif?

OpenStudy (amistre64):

i spose its all those backlogs from days gone by :) wonder if the sites fixed yet or if this is just a weekend break

myininaya (myininaya):

The site isn't completely fixed for me.

OpenStudy (anonymous):

not nearly

OpenStudy (amistre64):

emcormic had a meeting with preetha yesterday and were all going to be banned :)

myininaya (myininaya):

Sounds fascinating.

OpenStudy (anonymous):

good, them maybe i'll get some work done!

OpenStudy (amistre64):

work = force times distance ... so i dont see that being possible ;)

OpenStudy (anonymous):

more time to listen to spike jones, not to mention sleep goodnight friends

myininaya (myininaya):

It is still early.

OpenStudy (amistre64):

bon giorno :)

OpenStudy (anonymous):

:)

myininaya (myininaya):

You sleep as much as my cat. Goodnight.

OpenStudy (amistre64):

im glad i wasnt in THAT focus group

myininaya (myininaya):

We can find a different group for you later.

OpenStudy (amistre64):

... better be cupcakes ;)

myininaya (myininaya):

You can not belong to a set of cupcakes. But I suppose you could belong to a set of pretend cupcakes.

OpenStudy (amistre64):

like easy bake food preparations .... sounds fair to me

OpenStudy (freckles):

Moderators are weird.